The life of a 60- watt light bulb in hours is known to be normally distributed with σ = 25
hours. Create 5 different random samples of 100 bulbs each which has a mean life of x_bar ~
1000 hours and perform one-way ANOVA with state it.
1
Expert's answer
2020-08-17T19:20:46-0400

The total sample size is "N=500." Therefore, the total degrees of freedom are:


"df_{total}=500-1=499"

The between-groups degrees of freedom are "df_{between}=5-1=4," and the within-groups degrees of freedom are:


"df_{within}=df_{total}-df_{between}=499-4=495""\displaystyle\sum_{i,j}X_{ij}=499712""\displaystyle\sum_{i,j}X_{ij}^2=499691630""SS_{total}=\displaystyle\sum_{i,j}X_{ij}^2-\dfrac{1}{N}(\displaystyle\sum_{i,j}X_{ij})^2=267464.112""SS_{within}=\displaystyle\sum_{i}(n_i-1)s_i^2=266084.42""SS_{between}=\\displaystyle\\sum_{i}n_i(\\bar{x}-\\bar{\\bar{x}})^2=1379.692\n\u200b""MS_{between}=\dfrac{SS_{between}}{df_{between}}=\dfrac{1379.692}{4}=344.923""MS_{within}=\dfrac{SS_{within}}{df_{within}}=\dfrac{266084.42}{495}=537.544""F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{344.923}{537.544}=0.642"

The following null and alternative hypotheses need to be tested:

"H_0: \mu_1=\mu_2=\mu_3=\mu_4=\mu_5"

"H_1:" Not all means are equal.

The above hypotheses will be tested using an F-ratio for a One-Way ANOVA.


Based on the information provided, the significance level is "\alpha=0.05," and the degrees of freedom are "df_1=4" and "df_2=4," therefore, the rejection region for this F-test is "R=\{F:F>F_c=2.39\}."

Test Statistics


"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{344.923}{537.544}=0.642"

Since it is observed that "F=0.642<2.39=F_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 5 population means are equal, at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=0.633," and since "p=0.633\geq0.05,"

it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that not all 5 population means are equal, at the "\alpha=0.05" significance level.